# FedOC: Optimizing Global Prototypes with Orthogonality Constraints for Enhancing Embeddings Separation in Heterogeneous Federated Learning

### v1

@@ Abstract
Federated Learning (FL) has emerged as an essential framework for distributed machine learning, especially with its potential for privacy-preserving data processing. However, existing FL frameworks struggle to address statistical and model heterogeneity, which severely impacts model performance. While Heterogeneous Federated Learning (HtFL) introduces prototype-based strategies to address the challenges, current approaches face limitations in achieving optimal separation of prototypes. This paper presents FedOC, a novel HtFL algorithm designed to improve global prototype separation through orthogonality constraints, which not only increase intra-class prototype similarity but also significantly expand the inter-class angular separation. With the guidance of the global prototype, each client keeps its embeddings aligned with the corresponding prototype in the feature space, promoting directional independence that integrates seamlessly with the cross-entropy (CE) loss. We provide theoretical proof of FedOC’s convergence under non-convex conditions. Extensive experiments demonstrate that FedOC outperforms seven state-of-the-art baselines, achieving up to a 10.12% accuracy improvement in both statistical and model heterogeneity settings.
@@ 1 Introduction
In recent years, Federated Learning (FL) has emerged as a promising distributed machine learning paradigm [5]. FL eliminates the need for clients to exchange data, allowing data to remain decentralized, which makes it an favorable solution to data privacy challenges [1].
Previous work mainly focuses on homogeneous FL, which assumes that all client models are identical. However, in large-scale real-world scenarios, considerable disparities exist in both client data distribution, often termed statistical heterogeneity, and hardware resources, known as system heterogeneity [40, 16]. Data heterogeneity severely impacts the convergence and performance of global model [23]. Furthermore, discrepancies in hardware resources can result in model heterogeneity across clients, posing significant challenges to conventional FL approaches that rely on aggregating model parameters [40]. In addition, homogeneous FL trains a shared global model by exchanging gradients, which further imposes significant communication costs as well as privacy exposure risks [37, 43].
To tackle these challenges, Heterogeneous Federated Learning (HtFL) has emerged as a novel FL paradigm capable of handling both data heterogeneity and model heterogeneity simultaneously [33, 41, 43]. HtFL incorporates prototype-based learning, which communicates client prototypes rather than model gradients to the server, thus alleviating issues related to model diversity and communication costs. However, existing weighted-average HtFL solutions like FedProto [33] faces several limitations. First, aggregating global prototypes on the server side via weighted averaging requires clients to upload sample size, which may lead to leakage of data distribution information [41]. Second, due to statistical heterogeneity, the intra-class distribution discrepancy exists for the same class across clients. This disparity leads to inconsistent decision boundaries for the same class, which hinders the representation of a unified class distribution in the global feature space [43]. Furthermore, as presented in Fig. 1(a), the aggregated global prototypes suffer from limited separation margins, thereby diminishing their discriminative ability across classes.
@@
Figure 1: We train models on the CIFAR-10 dataset under a pathological partitioning scheme (each client is assigned two classes). We use t-SNE [36] to visualize their performance on previously unseen test data in the feature space. The results indicate that FedProto [33] exhibits weak feature separation, with the two prototypes nearly overlapping. FedTGP [43] increases prototype margin but still lacks sufficient distinction in the feature space. In contrast, our FedOC not only reduces inter-prototype similarity but also achieves high intra-class compactness, ensuring that feature representations of samples from the same class are tightly clustered. This outcome implies that FedOC enhances both classification and generalization abilities in settings with statistical and model heterogeneity.
@@
The recent HtFL solution FedTGP [43] has successfully improved separation by increases the Euclidean distance between prototypes. However, this approach still faces several limitations. First, contrastive learning, which works together with cross-entropy (CE) loss, primarily reduces the intra-class distance between prototypes but fails to explicitly enforce inter-class separation. Second, as shown in Figure 1(b), in high-dimensional spaces, the Euclidean distance makes it difficult to effectively distinguish different samples as distance variations between samples diminish [35]. Third, augmenting Euclidean distances fails to effectively leverage the angular separation characteristics inherent in the CE loss [26].
@@
To address these limitations, we propose a novel HtFL algorithm called FedOC, which optimizes global prototypes through orthogonality constraints (OC). FedOC improves intra-class prototype similarity while preserving semantic integrity and explicitly promotes angular separation between classes, ensuring that global prototypes achieve maximal directional independence in the feature space. Under the guidance of global prototypes, as demonstrated in Fig. 1(c), clients can ensure that the feature representations are directionally aligned with their corresponding global prototypes in the feature space, thereby facilitating a more effective integration with the angular properties inherent to CE loss.
@@
We evaluate FedOC against seven HtFL methods across three datasets with two statistical heterogeneity settings and multiple model heterogeneity scenarios. The experimental results show an improvement in accuracy by 10.12% over the best baseline under both statistical and model heterogeneity. Our contributions can be summarized as follow:
@@
  • We present a training framework FedOC for HtFL with orthogonality constraints. To the best of my knowledge, this is the first work to introduce orthogonality constraints into heterogeneous federated learning.
@@
  • We observed that increasing the Euclidean distance between prototypes is incapable of effectively integrating with the angular characteristics inherent to CE loss. Our proposed FedOC algorithm enhances intra-class clustering within the feature space while explicitly promoting angular separation between classes, which aligns with the CE loss.
@@
  • We offer a theoretical convergence guarantee for FedOC and rigorously establish the convergence rate under non-convex conditions.
@@ 2.1 Heterogenous Federated Learning
Federated Learning (FL) has gained significant attention as a privacy-preserving distributed machine learning approach, allowing collaborative model training without sharing private data [24]. The traditional FL methods, such as FedAvg [24], focus on training a global model across multiple clients by aggregating gradients [10]. However, these methods face major challenges in the presence of statistical and model heterogeneity, often resulting in suboptimal model performance due to non-IID data and diverse client hardware capabilities [15]. Although personalized FL methods [31, 17, 3] are introduced to address the statistical heterogeneity in FL, they remain unsuitable for scenarios where clients deploy heterogeneous models tailored to their specific tasks.
Heterogeneous Federated Learning (HtFL) has emerged as a solution that supports both model and statistical heterogeneity simultaneously, while ensuring user privacy. Early HtFL solutions [4, 8, 38] let clients sample different submodels from a shared global architecture, thereby reducing the computational burden on low-resource devices. However, these methods necessitate a common architecture to be shared across clients, raising concerns regarding the privacy of clients’ model architectures.
An alternative approach to system heterogeneity shares a common model of the same structure at the server side, e.g. a global header or a global generator, as a way to transfer global knowledge. LG-FedAvg [19] and FedGH [41] require clients to share the same top layer while allowing them to have different architectures in their lower layers. However, sharing and aggregation of top layers may result in suboptimal performance due to statistical heterogeneity [18, 22]. FedGen [44] is designed to mitigate model drift in heterogeneous FL environments by using a generative model to extract global knowledge from local models. While this method avoids reliance on proxy datasets, its performance highly relies on the quality of the generator [43]. Despite allowing some level of flexibility, they still assume that clients share partially similar models.
@@
Prototype-based FL, e.g. FedProto [33] and FedPCL [34], have demonstrated the potential in handling both model and statistical heterogeneity. These approaches not only improve model convergence in non-IID settings but also significantly reduce communication overhead by focusing on prototype updates instead of full model updates [34]. However, these methods rely on simple weighted averages of client prototypes, which increase the chance of overlap between global prototypes in the feature space. While FedTGP [43] increases prototype separation by contrastive learning without using weighted averages, boosting the Euclidean distance between embedding layers does not integrate well with CE loss [26, 35].
In this work, we apply orthogonality constraints to explicitly enlarge the angular margins between global prototypes. Under the guidance of the global prototypes, client embeddings maintain alignment with their corresponding global prototypes while preserving directional independence, facilitating a more effective integration with the angular characteristics inherent to CE loss.
@@ 2.2 Orthogonality
Orthogonality typically describes the directional independence between two vectors. In high-dimensional space, two vectors are considered orthogonal if their dot product is zero [26]. The previous formulations [13, 28] to achieve orthogonality in the feature space generally rely on computing singular value decomposition (SVD). However, these SVD-based methods are often computationally expensive. C-FSCIL [7], OPL [26] and POP [21] enhance feature separation between classes via orthogonality constraints. However, they are only applicable to the optimisation of centralised training scenarios and are not applicable to the optimisation of federated scenarios. FOT [2] and FedSOL [12] are primarily concerned with the orthogonality of update directions to minimise interference across clients or tasks. However, neither of them can guarantee the orthogonality of the samples in the feature space.
@@ 3.1 Problem Definition
We consider a problem involving $M$ clients, which have heterogeneous models and private data. The local model of client $k$, $k\in\{1,\dots,M\}$, is split into two components: a feature extractor $f_{k}$ parameterized by $\phi_{k}$, and a classifier $h_{k}$ parameterized by $\theta_{k}$. Given a sample pair ($x,y$), the feature extractor $f_{k}$ transforms $x\in\mathbb{R}^{D}$ into a feature representation $r\in\mathbb{R}^{K}$:
$r=f_{k}(x;\phi_{k}),$ (1)
where $D$ is the dimension of input and $K$ is the dimension of feature representation ($K\ll D$). The classifier $h_{k}$ further maps this $r$ to a probability distribution $output\in\mathbb{R}^{C}$:
$output=h_{k}(r;\theta_{k}),$ (2)
where $C$ is the total number of classes. The parameters of client model is defined as $\omega_{k}=(\phi_{k},\theta_{k})$.
The optimization objective of FedOC is defined as:
$\min\frac{1}{M}\sum_{k=1}^{M}\mathcal{L}_{k}(\mathcal{D}_{k},\omega_{k},\mathcal{P}),$ (3)
where $\mathcal{D}_{k}$ is the dataset of client $k$, $\mathcal{P}$ is the global prototype and $\mathcal{L}_{k}$ is the total loss of client $k$.
3.2 Orthogonality Constraints of Global Prototype
@@
To enhance the separation between global prototypes in the feature space, we adopt orthogonality constraints to train global prototypes. After client local training, the server will get the client prototype $P_{k}^{c}$ of class $c$, and prototypes of the remaining classes $P_{k}^{\bar{c}}$, where $\bar{c}\neq c$. Our purpose is to cluster $\hat{P}^{c}$ and $P_{k}^{c}$, while discriminating between $\hat{P}^{c}$ and $P_{k}^{\bar{c}}$. Formally, we define $s$ as intra-prototype similarity amd $d$ as inter-prototype similarity:
$s=\frac{1}{|B|}\sum_{P_{k}^{c}\in B}\langle P_{k}^{c},\hat{P}^{c}\rangle,$ (5)
$d=\frac{1}{|B|}\frac{1}{|C|}\sum_{P_{k}^{c}\in B}\sum_{c\neq\bar{c},\bar{c}\in C}|\langle P_{k}^{c},\hat{P}^{\bar{c}}\rangle|,$ (6)
@@
$\langle v_{i},v_{j}\rangle=\frac{v_{i}\cdot v_{j}}{\|v_{i}\|_{2}\cdot\|v_{j}\|_{2}},$ (7)
where $||\cdot||_{2}$ refers to the $l_{2}$ norm operator.
Then, we define a unified loss function $\mathcal{L}_{\mathrm{OC}}$ that simultaneously ensures intra-class clustering and inter-class orthogonality within a mini-batch as follows:
$\mathcal{L}_{\mathrm{OC}}=\lambda_{s}*(1-s)+\gamma*d,$ (8)
@@
The training objective is to minimize the loss function $\mathcal{L}_{\mathrm{OC}}$ defined in Eq. 8. Notice that $(1-s)>0$ inherently holds (as $s\in(-1,1)$, making $(1-s)\in(0,2)$), and $d$ is already an absolute value with $d\in(0,1)$. Therefore, we should maximize $s$ toward 1 while minimizing $d$ toward 0. As depicted in Fig. 2, when minimizing this overall loss, the first term $\lambda_{s}*(1-s)$ promotes clustering of prototypes within the same class, while the second term $\gamma*d$ enforces orthogonality among prototypes of different classes. The loss can be implemented efficiently in a vectorized manner on the mini-batch level, avoiding any loops.
@@
3.3 The Suitability of OC for CE Loss
The Cross Entropy (CE) loss function, defined as:
$\mathcal{L}_{\mathrm{CE}}(\hat{y},y)=-\log\frac{\exp{(r^{y}\cdot w^{y})}}{\sum_{j\neq y}\exp{(r^{y}\cdot w^{j})}},$ (9)
is widely used in deep learning models for classification tasks. We define the classifier $W=[w^{1},...,w^{C}]$, where $w^{y}\in\mathbb{R}^{K}$ is the learning projection vector of label $y$. The traditional CE loss can then be defined in terms of discrepancy between the predicted $\hat{y}$ and ground-truth label $y$, by projecting the features $r^{y}$ onto the weight matrix $W$. When a client trains the model with the global prototypes, it ensures that $r^{y}$ remains aligned with the prototype $\hat{P}^{y}$. During updates with SGD, the CE loss further reinforces the alignment between feature representation $r^{y}$ and its corresponding weight vector $w^{y}$ through a dot product operation, given as:

### v2

# FedORGP: Guiding Heterogeneous Federated Learning with Orthogonality Regularization on Global Prototypes

@@ Abstract
Federated Learning (FL) has emerged as an essential framework for distributed machine learning, especially with its potential for privacy-preserving data processing. However, existing FL frameworks struggle to address statistical and model heterogeneity, which severely impacts model performance. While Heterogeneous Federated Learning (HtFL) introduces prototype-based strategies to address the challenges, current approaches face limitations in achieving optimal separation of prototypes. This paper presents FedORGP, a novel HtFL algorithm designed to improve global prototype separation through orthogonality regularization, which not only encourages intra-class prototype similarity but also significantly expands the inter-class angular separation. With the guidance of the global prototype, each client keeps its embeddings aligned with the corresponding prototype in the feature space, promoting directional independence that integrates seamlessly with the cross-entropy (CE) loss. We provide theoretical proof of FedORGP’s convergence under non-convex conditions. Extensive experiments demonstrate that FedORGP outperforms seven state-of-the-art baselines, achieving up to 10.12% accuracy improvement in scenarios where statistical and model heterogeneity coexist.
@@ 1 Introduction
In recent years, Federated Learning (FL) has emerged as a promising distributed machine learning paradigm [7]. FL eliminates the need for clients to exchange data, allowing data to remain decentralized, which makes it an favorable solution to data privacy challenges [1, 25].
Previous work mainly focuses on homogeneous FL, which assumes that all client models are identical. However, in large-scale real-world scenarios, considerable disparities exist in both client data distribution, often termed statistical heterogeneity, and hardware resources, known as system heterogeneity [41, 18]. Data heterogeneity severely impacts the convergence and performance of global model [24, 17]. Furthermore, discrepancies in hardware resources can result in model heterogeneity across clients, posing significant challenges to conventional FL approaches that rely on aggregating model parameters [41, 13]. In addition, homogeneous FL trains a shared global model by exchanging gradients, which further imposes significant communication costs as well as privacy exposure risks [36, 44].
To tackle these challenges, Heterogeneous Federated Learning (HtFL) has emerged as a novel FL paradigm capable of handling both data heterogeneity and model heterogeneity simultaneously [32, 42, 44]. HtFL incorporates prototype-based learning, which communicates client prototypes rather than model gradients to the server, thus alleviating issues related to model diversity and communication costs. However, existing weighted average HtFL solutions like FedProto [32] faces several limitations. First, aggregating global prototypes on the server side via weighted averaging requires clients to upload sample sizes, which may lead to leakage of data distribution information [42]. Second, as shown in Fig.1(a), the weighted average prototype lacks a well-defined decision boundary, resulting in overlapping feature distributions among different classes in the feature space [44].
@@
Figure 1: We train models on the CIFAR-10 dataset and use t-SNE [35] to visualize their performance on previously unseen test samples (16 per class) within the feature space, with triangles indicating prototypes and circles denoting samples. The results indicate that FedProto [32] exhibits weak feature separation. FedTGP [44] increases prototype margin but still lacks sufficient distinction in the feature space. In contrast, our FedORGP can reduce inter-sample similarity and increase intra-class compactness, thereby effectively classifying the majority of samples. This result suggests that FedORGP exhibits robust generalization performance under conditions of both statistical and model heterogeneity.
@@
While the recent HtFL solution FedTGP [44] has successfully improved separation by increasing the Euclidean distance between prototypes, this approach still faces several limitations. First, contrastive learning, which works together with cross-entropy (CE) loss, primarily reduces the intra-class distance between prototypes but fails to explicitly enforce inter-class separation. Second, as shown in Fig.1(b), in high-dimensional spaces, the Euclidean distance makes it difficult to effectively distinguish different samples as distance variations between samples diminish [34]. Third, augmenting Euclidean distances fails to effectively leverage the angular separation characteristics inherent in CE loss [27].
@@
To address these limitations, we propose a novel HtFL algorithm called FedORGP, which optimizes global prototypes through orthogonality regularization. FedORGP improves intra-class prototype similarity while preserving semantic integrity and explicitly promotes angular separation between classes, ensuring that global prototypes achieve maximal directional independence in the feature space. Under the guidance of global prototypes, as demonstrated in Fig.1(c), clients can ensure that the feature representations are directionally aligned with their corresponding global prototypes in the feature space, thereby facilitating a more effective integration with the angular properties inherent to CE loss.
@@
We evaluate FedORGP against seven state-of-the-art HtFL methods across four datasets in scenarios where both data heterogeneity (practical distribution and Dirichlet distribution) and model heterogeneity (four levels of model heterogeneity) coexist. The experimental results show an improvement in accuracy by 10.12% over the best baseline under both statistical and model heterogeneity. Our contributions can be summarized as follows:
@@
  • We present a novel framework FedORGP for HtFL with orthogonality regularization. To the best of my knowledge, this is the first work to introduce orthogonality regularization into heterogeneous federated learning where data and model heterogeneity coexist.
@@
  • We observe that increasing the Euclidean distance between prototypes is incapable of effectively integrating with the angular characteristics inherent to CE loss. Our proposed FedORGP algorithm enhances intra-class clustering within the feature space while explicitly promoting angular separation between classes.
@@
  • We offer a theoretical convergence guarantee for FedORGP and rigorously establish the convergence rate under non-convex conditions.
@@ 2.1 Heterogenous Federated Learning
Heterogeneous Federated Learning (HtFL) has emerged as a solution that supports both model and statistical heterogeneity simultaneously, while ensuring user privacy. Early HtFL solutions [5, 10, 38] let clients sample different submodels from a shared global architecture, thereby reducing the computational burden on low-resource devices. However, these methods necessitate a common architecture to be shared across clients, raising privacy concerns regarding clients’ model architectures.
An alternative approach to system heterogeneity shares a common model of the same structure at the server side, e.g. a global header [42] or a global generator [45], as a way to transfer global knowledge. LG-FedAvg [20] and FedGH [42] require clients to share the same top layer while allowing them to have different architectures in their lower layers. However, sharing and aggregation of top layers may result in suboptimal performance due to statistical heterogeneity [19, 23]. FedGen [45] is designed to mitigate model drift in heterogeneous FL environments by using a generative model to extract global knowledge from local models. While this method avoids reliance on proxy datasets, its performance highly relies on the quality of the generator [44]. Despite allowing some level of flexibility, they still assume that clients share partially similar models.
@@
Prototype-based HtFL methods [32, 33, 40] have shown promise in addressing both model and statistical heterogeneity. These approaches not only enhance model convergence in non-IID settings but also substantially reduce communication overhead by transmitting prototypes instead of full models [33]. However, these methods rely on simple weighted averages of client prototypes, which increase the chance of overlap between global prototypes in the feature space. While FedTGP [44] increases prototype separation by contrastive learning without using weighted averages, boosting the Euclidean distance between embeddings does not integrate well with CE loss [27, 34]. In this work, we introduce orthogonal regularization to explicitly enhance the angular margin among global prototypes, thereby tackling data and model heterogeneity in HtFL.
2.2 Prototype Learning
Prototype refers to the average of features representations of the same class. FedNH [4] addresses data heterogeneity by leveraging uniformity and semantics of class prototypes. FPL [12] and FedPLVM [37] mainly focus on solving the domain drift problem using prototype clustering. However, they all require model aggregation, which necessitates that client models share the same architecture, implying that they cannot work in scenarios with model heterogeneity. Our proposed FedORGP encourages prototype separation to establish distinct margins among prototypes, thereby guiding heterogeneous client models to assimilate global knowledge.
@@ 2.3 Orthogonality
Orthogonality typically describes the directional independence between two vectors. In high-dimensional space, two vectors are considered orthogonal if their dot product is zero [27]. FediOS [6] effectively separates generic and personalized features using orthogonal projections to address feature skew in personalized federated learning. However, it requires all client models to have identical architectures. C-FSCIL [9], OPL [27] and POP [22] enhance feature separation between classes via orthogonality regularization. However, they are only applicable to the optimisation of centralised training scenarios and are not applicable to the optimisation of federated scenarios. FOT [2] and FedSOL [15] are primarily concerned with the orthogonality of update directions to minimise interference across clients or tasks. However, neither of them can guarantee the orthogonality of the samples in the feature space.
@@ 3.1 Problem Definition
We consider a problem involving $M$ clients, which have heterogeneous models and private data. The local model of client $k$, $k\in\{1,\dots,M\}$, is split into two components: a feature extractor $f_{k}$ parameterized by $\phi_{k}$, and a classifier $h_{k}$ parameterized by $\theta_{k}$. Given a sample pair ($x,y$), the feature extractor $f_{k}$ transforms $x\in\mathbb{R}^{D}$ into a feature representation $r=f_{k}(x;\phi_{k})$, $r\in\mathbb{R}^{K}$, where $D$ is the dimension of input and $K$ is the dimension of feature representation ($K\ll D$). The classifier $h_{k}$ maps the feature vector $r$ to logits $\in\mathbb{R}^{C}$, where $logits=h_{k}(r;\theta_{k})$ and $C$ denotes the total number of classes. The client model’s parameters are defined as $\omega_{k}=(\phi_{k},\theta_{k})$. The optimization objective of FedORGP is defined as:
$\min\sum_{k=1}^{M}\frac{|\mathcal{D}_{k}|}{|\mathcal{D}|}\mathcal{L}_{k}\left(\mathcal{D}_{k},\omega_{k},\mathcal{P}\right),$ (1)
where $|\mathcal{D}|$ denotes the total size of datasets across all clients, $|\mathcal{D}_{k}|$ indicates the size of the dataset of client $k$, $\mathcal{P}$ represents the global prototype.
3.2 Orthogonality Regularization on Prototypes
@@
To enhance the separation between global prototypes in the feature space, we adopt orthogonality regularization to train global prototypes. After client local training, the server will get the client prototypes $\mathcal{P}_{k}$, which consists of the prototypes of different classes of $k$-th client. Then, the dataset we use to train the global prototype can be expressed as $\mathcal{Q}=\bigcup_{k=1}^{M}\{P^{c}_{k}\mid c\in\mathcal{C}_{k}\}$, where $\mathcal{C}_{k}\subseteq\{0,1,...,C-1\}$ denotes the set of classes present on $k$-th client.
Our purpose is to cluster $\hat{P}^{c}$ and $P_{k}^{c}$, while discriminating between $\hat{P}^{c}$ and $P_{k}^{\bar{c}}$, where $\bar{c}\neq c$. Formally, within a mini-batch $B_{p}\subseteq\mathcal{Q}$, we define $s$ as intra-prototype similarity and $d$ as inter-prototype similarity:
$s$ $=\frac{1}{|B_{p}|}\sum_{P_{k}^{c}\in B_{p}}\langle P_{k}^{c},\hat{P}^{c}\rangle,$ (3)
$d$ $=\frac{1}{|B_{p}|}\frac{1}{|C|-1}\sum_{P_{k}^{c}\in B_{p}}\sum_{\begin{subarray}{c}\bar{c}\in C\\ c\neq\bar{c}\end{subarray}}\left|\langle P_{k}^{c},\hat{P}^{\bar{c}}\rangle\right|,$ (4)
@@
Then, we define a unified loss function $\mathcal{L}_{\mathrm{OR}}$ that simultaneously ensures intra-class clustering and inter-class orthogonality within a mini-batch as follows:
$\mathcal{L}_{\mathrm{OR}}=\lambda_{s}*(1-s)+\gamma*d,$ (5)
@@
The training objective is to minimize the loss function $\mathcal{L}_{\mathrm{OR}}$ defined in Eq. 5. Notice that $(1-s)>0$ inherently holds (as $s\in(-1,1)$, making $(1-s)\in(0,2)$), and $d$ is already an absolute value with $d\in(0,1)$. Therefore, we should maximize $s$ toward 1 while minimizing $d$ toward 0. As depicted in Fig.2, when minimizing this overall loss, the first term $\lambda_{s}*(1-s)$ promotes clustering of prototypes within the same class, while the second term $\gamma*d$ enforces orthogonality among prototypes of different classes. The loss can be implemented efficiently in a vectorized manner on the mini-batch level, avoiding any loops.
@@
3.3 The Suitability of Orthogonality for CE Loss
The multi-class Cross-Entropy loss function is a common loss function used in classification problems, for each sample $(x,y)$, the loss function can be defined as follows:
$\mathcal{L}_{CE}(\hat{y},y)$ $=-\sum_{c=1}^{C}\mathbb{I}(c=y)\log(p_{c})$
$=-\log\left(\frac{\exp(z_{y})}{\sum_{c^{\prime}=1}^{C}\exp(z_{c^{\prime}})}\right)$ (6)
$=-\log\left(\frac{\exp(r^{y}\cdot w^{y})}{\sum_{c^{\prime}=1}^{C}\exp(r^{y}\cdot w^{c^{\prime}})}\right)$
where $\hat{y}$ is the predicted label and $y$ is the truth label. $\mathbb{I}\left(\cdot\right)$ is an indicator function. $p_{c}$ represents the predicted probability that the model predicts the sample to belong to class $c$. $z_{y}$ is the logits of the sample. We define the classifier $W=[w^{1},...,w^{C}]$, where $w^{y}\in\mathbb{R}^{K}$ is the learning projection vector of label $y$. The CE loss can then be defined in terms of discrepancy between the predicted $\hat{y}$ and ground-truth label $y$, by projecting the features $r^{y}$ onto the weight matrix $W$. When a client trains the model with the global prototypes, it ensures that $r^{y}$ remains aligned with the prototype $\hat{P}^{y}$. During updates with SGD, the CE loss further reinforces the alignment between weight vector $w^{y}$ and its corresponding feature representation $r^{y}$ through a dot product operation, given as: